# Proton-air interactions at ultra-high energies in muon-depleted air showers with different depths

### v1

@@ Abstract
In this work, we find that this steepness, quantified by the parameter $\Lambda_{\mu}$, evolves with the depth of the shower maximum, $X_{\max}$, assuming a universal value for shallow showers and an enhanced dependence on the high-energy hadronic interaction model for deep showers. We show that $X_{\max}$ probes the so-called hadronic activity of the first interaction, thus allowing direct access to the energy spectrum of neutral pions in different regions of the kinematic phase space of the first interaction.
@@ 1 Introduction
Recently, it was shown that the fluctuations in the muon content of extensive air showers (EAS) are sensitive to fluctuations in the energy sharing among secondaries of the first $p$-air interaction [2]. Additionally, the distribution of the number of muons in proton-induced EAS has a so-called tail towards low values, constituted by muon-depleted showers. The steepness of this tail, measured by $\Lambda_{\mu}$, probes the production cross-section of neutral pions in the forward region of the kinematic phase-space of $p$-air interactions [3]. This measurement is resilient to different composition scenarios.
@@ 2 Evolution of $\Lambda_{\mu}$ with $X_{\rm max}$
In proton-induced extensive air showers, the fluctuations of the depth of the shower maximum, $X_{\max}$, and those of the number of muons reaching the ground level, $N_{\mu}$, are coupled, as both observables are sensitive to the multiparticle production properties of high-energy hadronic interactions. This coupling is clear from the joint distribution function $f$($X_{\rm max}$, $\ln N_{\mu}$), shown in the lower panel of Fig. 1. This figure was produced using $10^{6}$ proton-induced showers simulated with Conex v7.50 [4, 5] with primary energy $E_{0}=10^{19}\,$eV and zenith angle $\theta=67^{\circ}$, using the high-energy hadronic interaction model EPOS-LHC[6]. Inclined showers were chosen to mitigate artificial correlations between $X_{\rm max}$ and $\ln N_{\mu}$ through the truncation of the shower development. ***Furthermore, particles above $E_{\text{th}}=0.005\times E_{0}$ and their interactions are tracked individually, while below it, the longitudinal shower profile numerically solves cascade equations. The ground level was set to 1 40014001\,400\,1 400m a.s.l, the average height of the Pierre Auger Observatory [7], corresponding to an average vertical depth of $X_{\text{gr}}=880\,\mathrm{g\,cm^{-2}}$. The number of muons at the ground level, $N_{\mu}$, is defined by the value of the muon longitudinal profile at a slanted depth of $X=X_{\text{gr}}\sec\theta$. Only muons with $E_{\mu}>1\,$GeV were considered. The value of $X_{\rm max}$ is taken from a Gaisser-Hillas fit to the longitudinal profile of all charged particles.
@@
Figure 1: Lower panel: Joint distribution of ($X_{\rm max}$, $\ln N_{\mu}$), $f$($X_{\rm max}$, $\ln N_{\mu}$). The white contours contain $68\%$ (dotted), $95\%$ (dashed) and $99.7\%$ (solid) of the events. The horizontal lines represent bins in $X_{\rm max}$ with edges {650, 750, 800, 850, 900, $\{650,\,750,\,800,\,850,\,900,\,1300\}\mathrm{\,g\,cm^{-2}}$ and centers indicated by the triangular markers. Higher values of $X_{\rm max}$ are represented by darker shades of blue and lower ones by light shades of green. Upper panel: distribution of $\ln N_{\mu}$ for each bin in $X_{\rm max}$, $f_{X}(\ln N_{\mu})$, matching the colour of the corresponding $X_{\rm max}$ bin. Details of the simulations are written in the text.
@@
Preliminary studies have partially exploited $f$($X_{\rm max}$, $\ln N_{\mu}$) to constrain the macroscopic multi-particle production properties of the highest energy hadronic-interactions [8]. Although this macroscopic picture is useful, it does not allow constraining more fundamental quantities, such as the cross-section for the production of neutral pions, which certainly impacts the development of the electromagnetic cascade through their decay into photons. Furthermore, the correlation between $X_{\max}$ and $N_{\mu}$ is decreased by the fluctuations of the depth of the $p$-air interaction, $X_{0}$, which are independent of muon production. However, since $X_{0}$ cannot be experimentally accessed, we carry our investigations using $X_{\rm max}$ directly.
@@
The evolution of the steepness of the tail of the distribution of $\ln N_{\mu}$ with $X_{\rm max}$ is quantified through the change in the parameter $\Lambda_{\mu}$. The value of $\Lambda_{\mu}$ is extracted from unbinned likelihood fit to $f_{X}(\ln N_{\mu})$ assuming the function $y=A_{\mu}\exp\left\{\ln N_{\mu}/\Lambda_{\mu}\right\}$. Therefore, smaller (larger) values of $\Lambda_{\mu}$ correspond to a steeper (flatter) tail. The tail is defined by all values of $\ln N_{\mu}$ below the $10\%$ quantile of $f_{X}(\ln N_{\mu})$. We verified that the value of $\Lambda_{\mu}$ is independent of the choice of the quantile defining the tail provided it is below the bulk of the distribution. The values of $\Lambda_{\mu}$ in each bin of $X_{\rm max}$ is shown in Figure 2 for three different high-energy hadronic interaction models, EPOS-LHC  (blue), QGSJet-II.04 [11] (orange) and Sibyll 2.3d [12] (purple). The insets of the same figure display the distributions of $\ln N_{\mu}$ for the shallowest (left inset) and deepest (right inset) bins of $X_{\rm max}$.
Figure 2: $\Lambda_{\mu}$ as a function of $X_{\rm max}$ for the hadronic interaction models EPOS-LHC , QGSJet-II.04 and Sibyll 2.3d, for proton-induced EAS with $E_{0}=10^{19}$ eV and $\theta=67^{\circ}$. The distributions of $N_{\mu}$ for the shallowest and deepest bins in $X_{\rm max}$ are shown in the left and right insets, respectively.
Evidently, there is a monotonic increase of $\Lambda_{\mu}$ with $X_{\rm max}$, irrespective of the high-energy hadronic interaction model. Furthermore, the predicted value of $\Lambda_{\mu}$ is mildly model-independent for shallower showers and strongly model-dependent in deeper showers, where the relative model difference reaches $30\%$. For reference, there is a model dependence of $20\%$ in $\Lambda_{\mu}$ if extracted by fitting the distribution of the number of muons integrated over all $X_{\rm max}$.
@@ 3 Connection with the first interaction
Lastly, from the right panel of Figure 3, it is evident that deeper showers tend to correspond to first interactions with fast neutral pions. These immediately decay into high-energy photons that dominate the electromagnetic profile and yield deeper showers. So, the energy spectrum of neutral pions constrained to deep showers is harder, reducing the steepness tail of the distribution of $\ln N_{\mu}$. Therefore, the production spectrum of neutral pions can be constrained, as a function of the hadronic activity of the first $p$-air interaction, through the evolution of $\Lambda_{\mu}$ with $X_{\rm max}$, in phase-space regions unconstrained by accelerator data.
Finally, we propose a possible explanation for the increasing model dependence of $\Lambda_{\mu}$ with $X_{\rm max}$. Shallow showers are characterized by high hadronic activity (high multiplicity of secondaries among which energy is more evenly distributed), so secondaries are produced in kinematic regions more accessible to accelerator experiments. Thus the shape of the energy spectrum of secondary neutral pions is better constrained by data, yielding similar values of $\Lambda_{\mu}$ across models. In contrast, deep showers have reduced hadronic activity (low multiplicity of secondaries among which energy is very asymmetrically distributed). Thus, leading particles tend to occupy the forward region of the p-air interaction’s kinematic phase space, where the models are less constrained by accelerator data. Due to this freedom, the shape of the energy spectrum of neutral pions becomes more on the particular physical mechanisms employed by each model, resulting in an enhanced model dependence of $\Lambda_{\mu}$.
@@ 4 Experimental feasibility of the measurement of $\Lambda_{\mu}$ as a function of $X_{\rm max}$
The feasibility of the measurement of $\Lambda_{\mu}$ in each bin of $X_{\rm max}$ was quantified for two different composition scenarios. For primary energies in $[10^{18},10^{18.5}]$ eV, these scenarios correspond to the ones with the lowest and highest proton fractions given by the 4-mass fraction fit to Auger data presented in [18]. Each scenario is mimicked by sampling from $10^{5}$ proton, helium, nitrogen and iron-induced simulated showers, in proportions: p : He : N : Fe $=7:1:2:0$ and $1:3:1:0$. Furthermore, Gaussian resolutions with $\sigma(N_{\mu})/N_{\mu}=20\%$ [19] and $\sigma(X_{\max})=20\mathrm{\,g\,cm^{-2}}$ [20] were considered in the reconstruction of $\ln N_{\mu}$ and $X_{\rm max}$, respectively. Note that, in a real scenario, the composition interpretation of the cosmic ray flux depends on the hadronic interaction model and evolves with the primary energy [21]. Moreover, the falling cosmic ray spectrum in interplay with the composition couples the composition evolution with the precision of the measurement of $\Lambda_{\mu}$ via the limiting statistics. For simplicity, we decouple these effects by studying extreme composition cases in a small lower energy bin.
Figure 4: Fitted $\ln N_{\mu}$ distributions for each $X_{\rm max}$ bin for a mixture of p : He : N : Fe in proportion 1 : 3 : 1 : 0, assuming Gaussian resolutions of $20\%$ in $N_{\mu}$ and $20\,\mathrm{g\,cm^{-2}}$ in $X_{\max}$. All $10^{5}$ Conex showers were simulated with a primary energy of $E_{0}=10^{19}\,$eV and $\theta=67^{\circ}$ using the high-energy hadronic interaction model EPOS-LHC .
@@
We quantify the impacts of the reconstruction resolution and the presence of heavier primaries on $\Lambda_{\mu}$ separately. That is, for the composition scenario $1:3:1:0$, we fit the tail of each distribution of $\ln N_{\mu}$ to extract $\Lambda_{\mu}^{\mathrm{rec}}$, for two cases: a perfect reconstruction resolution in $N_{\mu}$ and $\sigma(N_{\mu})/N_{\mu}=20\%$. The resolution in $X_{\rm max}$ can be regarded as an additional smearing in $N_{\mu}$, so here we set it to $\sigma(X_{\max})=0\,\mathrm{g\,cm^{-2}}$. For these two cases, we compute the relative residuals between $\Lambda_{\mu}^{\text{rec}}$ and $\Lambda_{\mu}^{p}$, the value obtained for proton-induced showers, as shown in Figure 5, as a function of $X_{\rm max}$, using the three high-energy hadronic interaction model EPOS-LHC , QGSJet-II.04 and Sibyll 2.3d. Dashed lines correspond to a mixed composition with perfect resolution, and the solid lines to an assumed resolution of $20\%$ in the reconstruction of $N_{\mu}$. The black dotted line represents the model-averaged bias, and the grey line represents the difference between high-energy hadronic interaction models. We take this dependence as a systematic uncertainty.
The bias induced by a mixture of primary masses on $\Lambda_{\mu}$ is less than $5\%$, irrespective of the $X_{\rm max}$ bin. The small bias can be explained by the superposition principle: for a primary with a mass number $A$, the distribution of the number of muons is approximately the $A$-fold self-convolution of the one present for proton showers, thus suppressing the distribution’s tail. Hence, the tail of the distribution of $\ln N_{\mu}$ for all showers is dominated by the proton-induced ones, extending the results of Section 2 to realistic compositions. On the other hand, the convolution of the detector resolution with the distribution of $\ln N_{\mu}$ greatly biases $\Lambda_{\mu}$ in shallower showers since the tail of the distribution is very steep. Additionally, for the shallow showers, the bias is model-dependent. For deeper showers, the bias in $\Lambda_{\mu}$ reduces to about $10\%$, allowing for an accurate estimation of $\Lambda_{\mu}$. Detailed strategies to de-convolve the detector resolution from the physical fluctuations of $N_{\mu}$ could improve the bias in $\Lambda_{\mu}$ for shallower showers. However, for the sake of simplicity, we apply an average bias correction to $\Lambda_{\mu}$ and take the model dependence as a systematic uncertainty. This uncertainty in $\Lambda_{\mu}$ is greater than the difference between hadronic interaction models for showers with $X_{\max}<800\,\mathrm{g\,cm^{-2}}$, thus preventing model discrimination. Nevertheless, a measurement of $\Lambda_{\mu}$ for shallow showers could still allow the simultaneous exclusion of all models. Finally, these observations also apply to the scenario with the higher fraction of proton primaries.
@@
For showers with $X_{\max}>800\,\mathrm{g\,cm^{-2}}$, we determined the minimum number of events, $n_{\min}$, necessary to distinguish between $\Lambda_{\mu}$ predictions from different hadronic interaction models. The distinction threshold was set to $1$ and $3\,\sigma$, corresponding to $n_{\min}^{1\sigma}$ and $n_{\min}^{3\sigma}$ events, respectively. The error in $\Lambda_{\mu}$ was computed via a bootstrapping method to ensure its accurate estimation for low statistics. The values of $n_{\min}$ for the proton-rich and proton-poor scenarios can be found in Table 1. We assumed the aforementioned reconstruction resolutions.
@@
Since the models are close to indistinguishable for showers in the bulk of the $X_{\rm max}$ distribution, the number of events needed for a $3\sigma$ model separation is greater than the number of events in the samples. However, about 3 00030003\,0003 000 events allow a $1\sigma$ distinction between models in the deepest $X_{\max}$ bin, even in the proton-poor scenario. A $9$-fold increase in statistics allows for a $3\sigma$ distinction. Note that the acceptance and quality cuts applied to data from different experiments might require more events than those estimated here. Nevertheless, these numbers of events are within the reach of the Pierre Auger Observatory [22, 23]. For this experiment, the shower-to-shower estimations of $N_{\mu}$ and $X_{\rm max}$ could be provided by neural networks [24, 25, 26], shower universality [27] and/or the improved detectors of AugerPrime [28].

### v2

@@ Abstract
In this work, we find that this steepness, quantified by the parameter $\Lambda_{\mu}$, evolves with the depth of the shower maximum, $X_{\max}$, assuming a universal value for shallow showers and an enhanced dependence on the high-energy hadronic interaction model for deep showers. We show that $X_{\max}$ probes the so-called hadronic activity of the first proton-air interaction, thus allowing direct access to the energy spectrum of neutral pions in different regions of its kinematic phase space.
@@ 1 Introduction
Recently, it was shown that the fluctuations in the muon content of extensive air showers (EAS) are sensitive to fluctuations in the energy sharing among secondaries of the first $p$-air interaction [2]. Additionally, the distribution of the number of muons in proton-induced EAS has a so-called tail towards low values, constituted by muon-depleted showers. The steepness of this tail, measured by $\Lambda_{\mu}$, probes the production cross-section of neutral pions in the forward region of the kinematic phase-space of proton-air interactions [3]. This measurement is resilient to different composition scenarios.
@@ 2 Evolution of $\Lambda_{\mu}$ with $X_{\rm max}$
In proton-induced extensive air showers, the fluctuations of the depth of the shower maximum, $X_{\max}$, and those of the number of muons reaching the ground level, $N_{\mu}$, are coupled, as both observables are sensitive to the multiparticle production properties of high-energy hadronic interactions. This coupling is clear from the joint distribution function $f$($X_{\rm max}$, $\ln N_{\mu}$), shown in the lower panel of Figure 1. This figure was produced using $10^{6}$ proton-induced showers simulated with Conex v7.50 [4, 5] with primary energy $E_{0}=10^{19}\,$eV and zenith angle $\theta=67^{\circ}$, using the high-energy hadronic interaction model EPOS-LHC  [6]. Inclined showers were chosen to mitigate artificial correlations between $X_{\rm max}$ and $\ln N_{\mu}$ through the truncation of the shower development. ***Furthermore, particles above $E_{\text{th}}=0.005\times E_{0}$ and their interactions are tracked individually, while below it, the longitudinal shower profile numerically solves cascade equations. The ground level was set to 1 40014001\,400\,1 400m a.s.l, the average height of the Pierre Auger Observatory [7], corresponding to an average vertical depth of $X_{\text{gr}}=880\,\mathrm{g\,cm^{-2}}$. The number of muons at the ground level, $N_{\mu}$, is defined by the value of the muon longitudinal profile at a slanted depth of $X=X_{\text{gr}}\sec\theta$. Only muons with $E_{\mu}>1\,$GeV were considered. The value of $X_{\rm max}$ is taken from a Gaisser-Hillas fit to the longitudinal profile of all charged particles.
@@
Figure 1: Lower panel: Joint distribution of ($X_{\rm max}$, $\ln N_{\mu}$), $f$($X_{\rm max}$, $\ln N_{\mu}$). The white contours contain $68\%$ (dotted), $95\%$ (dashed) and $99.7\%$ (solid) of the shower simulations. The horizontal lines represent bins in $X_{\rm max}$ with edges {650, 750, 800, 850, 900, $\{650,\,750,\,800,\,850,\,900,\,1300\}\mathrm{\,g\,cm^{-2}}$ and centers indicated by the triangular markers. Higher values of $X_{\rm max}$ are represented by darker shades of blue and lower ones by light shades of green. Upper panel: distribution of $\ln N_{\mu}$ for each bin in $X_{\rm max}$, $f_{X}(\ln N_{\mu})$, matching the colour of the corresponding $X_{\rm max}$ bin. Details of the simulations are given in the main text.
@@
Preliminary studies have partially exploited $f$($X_{\rm max}$, $\ln N_{\mu}$) to constrain the macroscopic multi-particle production properties of the highest energy hadronic-interactions [8]. Although this macroscopic picture is useful, it does not allow the probing of more fundamental quantities, such as the cross-section for the production of neutral pions, which certainly impacts the development of the electromagnetic cascade through their decay into photons. Furthermore, the correlation between $X_{\max}$ and $N_{\mu}$ is decreased by the fluctuations of the depth of the $p$-air interaction, $X_{0}$, which are independent of muon production. However, since $X_{0}$ cannot be experimentally accessed, we carry our investigations using $X_{\rm max}$ directly.
@@
The evolution of the steepness of the tail of the distribution of $\ln N_{\mu}$ with $X_{\rm max}$ is quantified through the change in the parameter $\Lambda_{\mu}$, defined in the previous paragraph. Note that smaller (larger) values of $\Lambda_{\mu}$ correspond to a steeper (flatter) tail. We have verified that the value of $\Lambda_{\mu}$ is independent of the choice of the quantile defining the tail provided it is below the bulk of the distribution. The values of $\Lambda_{\mu}$ in each bin of $X_{\rm max}$ are shown in Figure 2 for three different high-energy hadronic interaction models, EPOS-LHC  (blue), QGSJet-II.04 [12] (orange) and Sibyll 2.3d [13] (purple). The insets of the same figure display the distributions of $\ln N_{\mu}$ for the shallowest (left inset) and deepest (right inset) bins of $X_{\rm max}$.
Figure 2: $\Lambda_{\mu}$ as a function of $X_{\rm max}$ for the hadronic interaction models EPOS-LHC , QGSJet-II.04 and Sibyll 2.3d, for proton-induced EAS with $E_{0}=10^{19}$ eV and $\theta=67^{\circ}$. The distributions of $\ln N_{\mu}$ for the shallowest and deepest bins in $X_{\rm max}$ are shown in the left and right insets, respectively.
Evidently, there is a monotonic increase of $\Lambda_{\mu}$ with $X_{\rm max}$, irrespective of the high-energy hadronic interaction model. Furthermore, the predicted value of $\Lambda_{\mu}$ is mildly model-independent for shallower showers and strongly model-dependent in deeper showers, where the relative model difference reaches $30\%$. For reference, there is a model dependence of $20\%$ in $\Lambda_{\mu}$ if extracted by fitting the distribution of the number of muons integrated over all $X_{\rm max}$. The described behaviours are qualitatively independent of the widths of the bins in $X_{\max}$, provided the same region of the $X_{\max}$ distribution is being probed.
@@ 3 Connection with the first interaction
Lastly, from the right panel of Figure 3, it is evident that deeper showers tend to correspond to first interactions with fast neutral pions. These immediately decay into high-energy photons that dominate the electromagnetic profile and yield deeper showers. So, the energy spectrum of neutral pions constrained to deep showers is harder, flattening the tail of the distribution of $\ln N_{\mu}$. Therefore, the production spectrum of neutral pions can be constrained, as a function of the hadronic activity of the first $p$-air interaction, through the evolution of $\Lambda_{\mu}$ with $X_{\rm max}$, in phase-space regions unconstrained by accelerator data. Note, however, that the divergence of the value of $\Lambda_{\mu}$ obtained with EPOS-LHC , relative to other models, in deeper bins in $X_{\max}$ cannot be solely explained from the differences in the spectrum of neutral pions of the first proton-air interaction.
Finally, we propose a possible explanation for the increasing model dependence of $\Lambda_{\mu}$ with $X_{\rm max}$. Shallow showers are characterized by high hadronic activity (high multiplicity of secondaries among which energy is more evenly distributed), so secondaries are produced in kinematic regions more accessible to accelerator experiments. Thus the shape of the energy spectrum of secondary neutral pions is better constrained by data, yielding similar values of $\Lambda_{\mu}$ across models. In contrast, deep showers have reduced hadronic activity (low multiplicity of secondaries among which energy is very asymmetrically distributed). Thus, leading particles tend to occupy the forward region of the p-air interaction’s kinematic phase space, where the models are less constrained by accelerator data. Due to this freedom, the shape of the energy spectrum of neutral pions becomes more dependent on the particular physical mechanisms employed by each model, resulting in an enhanced model dependence of $\Lambda_{\mu}$.
@@ 4 Experimental feasibility of the measurement of $\Lambda_{\mu}$ as a function of $X_{\rm max}$
The feasibility of the measurement of $\Lambda_{\mu}$ in each bin of $X_{\rm max}$ was quantified for two different composition scenarios. For primary energies in $[10^{18},10^{18.5}]$ eV, these scenarios correspond to the ones with the lowest and highest proton fractions given by the 4-mass fraction fit to Auger data presented in [19]. Each scenario is mimicked by sampling from $10^{5}$ proton, helium, nitrogen and iron-induced simulated showers, in proportions: p : He : N : Fe $=7:1:2:0$ and $1:3:1:0$. Furthermore, Gaussian resolutions with $\sigma(N_{\mu})/N_{\mu}=20\%$ [20] and $\sigma(X_{\max})=20\mathrm{\,g\,cm^{-2}}$ [11] were considered in the reconstruction of $\ln N_{\mu}$ and $X_{\rm max}$, respectively. In a real scenario, the composition interpretation of the cosmic ray flux depends on the hadronic interaction model and evolves with the primary energy [21]. Moreover, the falling cosmic ray spectrum in interplay with the composition couples the composition evolution with the precision of the measurement of $\Lambda_{\mu}$ via the limiting statistics. For simplicity, we decouple these effects by studying extreme composition cases in a small bin of primary energies.
Figure 4: Fitted distributions of $\ln N_{\mu}$ for each $X_{\rm max}$ bin for a mixture of p : He : N : Fe in proportion 1 : 3 : 1 : 0, assuming Gaussian resolutions of $20\%$ in $N_{\mu}$ and $20\,\mathrm{g\,cm^{-2}}$ in $X_{\max}$. The $10^{5}$ Conex showers, for each primary, were simulated with a primary energy of $E_{0}=10^{19}\,$eV and $\theta=67^{\circ}$ using the high-energy hadronic interaction model EPOS-LHC .
@@
We quantify the impacts of the reconstruction resolution and the presence of heavier primaries on $\Lambda_{\mu}$ separately. That is, for the composition scenario $1:3:1:0$, we fit the tail of each distribution of $\ln N_{\mu}$ to extract the reconstructed value of $\Lambda_{\mu}$, $\Lambda_{\mu}^{\text{rec}}$, for two cases: a perfect reconstruction resolution in $N_{\mu}$ and $\sigma(N_{\mu})/N_{\mu}=20\%$. The resolution in $X_{\rm max}$ can be regarded as an additional smearing in $N_{\mu}$, so here we set it to $\sigma(X_{\max})=0\,\mathrm{g\,cm^{-2}}$. For these two cases, we compute the residuals of $\Lambda_{\mu}^{\text{rec}}$ relative to the value of $\Lambda_{\mu}$ obtained for proton-induced showers, henceforth $\Lambda_{\mu}^{p}$, as shown in Figure 5, as a function of $X_{\rm max}$, using the three high-energy hadronic interaction model EPOS-LHC , QGSJet-II.04 and Sibyll 2.3d. Dashed lines correspond to a mixed composition with perfect resolution, and the solid lines to an assumed resolution of $20\%$ in the reconstruction of $N_{\mu}$. The black dotted line represents the model-averaged bias, and the grey line represents the difference between high-energy hadronic interaction models. We take this dependence as a systematic uncertainty.
The bias induced by a mixture of primary masses on $\Lambda_{\mu}$ is less than $5\%$, irrespective of the $X_{\rm max}$ bin. The small bias can be explained by the superposition principle: for a primary with a mass number $A$, the distribution of the number of muons is approximately the $A$-fold self-convolution of the one present for proton showers, thus suppressing the distribution’s tail. Hence, the tail of the distribution of $\ln N_{\mu}$ for all showers is dominated by the proton-induced ones, ensuring the applicability of the results of Section 2 to realistic compositions. On the other hand, the convolution of the detector resolution with the distribution of $\ln N_{\mu}$ greatly biases $\Lambda_{\mu}$ in shallower showers since the tail of the distribution is very steep. Additionally, for the shallow showers, the bias in $\Lambda_{\mu}$ is model-dependent. For deeper showers, the bias in $\Lambda_{\mu}$ reduces to about $10\%$, allowing for an accurate estimation of $\Lambda_{\mu}$. Detailed strategies to de-convolve the detector resolution from the physical fluctuations of $N_{\mu}$ could mitigate the bias in $\Lambda_{\mu}$ for shallower showers. However, for the sake of simplicity, we apply an average bias correction to $\Lambda_{\mu}$ and take the model dependence as a systematic uncertainty. This uncertainty in $\Lambda_{\mu}$ is greater than the difference between hadronic interaction models for showers with $X_{\max}<800\,\mathrm{g\,cm^{-2}}$, thus preventing model discrimination. Nevertheless, a measurement of $\Lambda_{\mu}$ for shallow showers could still allow the simultaneous exclusion of all models. Finally, these observations also apply to the scenario with the higher fraction of proton primaries.
@@
For showers with $X_{\max}>800\,\mathrm{g\,cm^{-2}}$, we determined the minimum number of events, $n_{\min}$, necessary to distinguish between $\Lambda_{\mu}$ predictions from different hadronic interaction models. The distinction threshold was set to $1$ and $3\,\sigma$, corresponding to $n_{\min}^{1\sigma}$ and $n_{\min}^{3\sigma}$ events, respectively. The error in $\Lambda_{\mu}$ was computed via a bootstrapping method to ensure its accurate estimation for low statistics. The values of $n_{\min}$ for the proton-rich and proton-depleted composition scenarios can be found in Table 1. We assumed the aforementioned reconstruction resolutions.
@@
Since the models are close to indistinguishable for showers in the bulk of the $X_{\rm max}$ distribution, the number of events needed for a $3\sigma$ model separation is greater than the number of events in the samples. However, about 3 00030003\,0003 000 events allow a $1\sigma$ distinction between models in the deepest $X_{\max}$ bin, even in the proton-poor scenario. A $9$-fold increase in statistics allows for a $3\sigma$ distinction. Note that the acceptance and quality cuts applied to data from different experiments might require more events than those estimated here. Nevertheless, these numbers of events are within the reach of the Pierre Auger Observatory [22, 23]. For this experiment, the shower-to-shower estimations of $N_{\mu}$ and $X_{\rm max}$ could be provided by neural networks [24, 25, 26, 27], shower universality [28] and/or the improved detectors of AugerPrime [29].